# Network Analytics for Anti-Money Laundering – A Systematic Literature Review and Experimental Evaluation

### v2

@@ 3.3 Top Cited Papers
A final step in our discussion is an in-depth analysis of the top-cited papers. These papers are well received by the scientific community, meaning that the methods that are applied in these papers are likely to be of high quality. To this end, we select the $10\%$ most cited papers per year, and complement these with the $10\%$ most cited papers overall (see Section 2.1). This results in 20 papers as enlisted in Table 3.
Reference
Title
@@
[91] The application of data mining techniques in financial fraud detection: A classification framework and an academic review of literature 2011 1307
@@
[38] Using social network analysis to prevent money laundering 2017 200
@@
[23] WireVis: Visualization of Categorical, Time-Varying Data From Financial Transactions 2007 171
@@
[105] Financial Crimes Enforcement Network AI System (FAIS) Identifying Potential Money Laundering from Reports of Large Cash Transactions 1995 152
@@
[31] The application of social network analysis algorithms in a system supporting money laundering detection 2015 141
[80] Visualizing Dynamic Bitcoin Transaction Patterns 2016 122
@@
[39] A framework for data mining-based anti-money laundering research 2007 106
@@
[125] Detecting Mixing Services via Mining Bitcoin Transaction Network With Hybrid Motifs 2021 94
@@
[79] Using friends for money: the positional importance of money-launderers in organized crime 2013 66
[139] A Local Algorithm for Structure-Preserving Graph Cut 2017 63
[72] FlowScope: Spotting Money Laundering Based on Graphs 2020 54
@@
[59] Illicit Bitcoin transactions: challenges in getting to the who, what, when and where 2018 47
@@
[97] EGRET: Extortion Graph Exploration Techniques in the Bitcoin Network 2018 38
@@
[69] Deep Learning and Explainable Artificial Intelligence Techniques Applied for Detecting Money Laundering–A Critical Review 2021 38
@@
[138] No Smurfs: Revealing Fraud Chains in Mobile Money Transfers 2014 32
@@
[15] Mapping the underground: Supervised discovery of cybercrime supply chains 2019 24
[93] Covertness Centrality in Networks 2012 20
@@
[24] Event-based approach to money laundering data analysis and visualization 2010 12
@@
[44] The effect of anti-money laundering policies: an empirical network analysis 2022 11
@@
[140] Behavior-Aware Account De-Anonymization on Ethereum Interaction Graph 2022 7
Table 3: Top-cited papers. Summary of the top-cited papers, with indication of whether it is a review and/or crypto-related.
@@
The results show that five out of 11 review papers are selected. This is to be expected, since review papers have a higher likelihood of being cited [30]. Additionally, five papers deal with crypto-related money laundering, indicating the strong research interest in these payment methods. Two causes that explain the growing interest are mentioned in literature. On the one hand, crypto currencies are widely used by criminals because of the absence of regulation [37] and the anonymity it provides [59, 122]. On the other hand, a full history of ownership and transactions is available on the blockchain for each coin, facilitating the discovery of transaction flows and patterns in payment networks [59, 122].
@@
The papers are classified according to the framework presented in [17] based on the processing steps and challenges presented. Table 4 summarises the papers over the different processing steps. The most prevalent pre-processing steps are feature engineering and exploration. The feature engineering is done to enhance GNN models and define different risk profiles. We see in the literature that this is either based on neighbourhood and centrality metrics [38, 31, 93] or on the aggregation of local (node-specific) information [140, 138]. When looking at exploration, some of the research is done on how to construct meaningful visualisations [23, 80], while other papers rely on clustering the nodes as additional support for their findings [79, 38].
@@
For the post-processing step, most research covers statistical evaluation, implementation and interpretation. The statistical evaluation is used to have a quantitative evaluation of the performance of the method. All research on the implementation looks at the efficiency of the methods. It deals solely with computation time and scalability. To have interpretations in the post-processing step, all papers mentioned use specific interpretable features.
@@
Pre-processing Sampling [38]
Exploration [38],[23],[80], [79]
Missing Value Treatment [38]
@@
Categorization, standardization & segmentation [31], [15]
Feature Engineering [38], [31], [93],[140], [138]
@@
Processing Unsupervised Learning [23], [105], [31], [79], [139], [72], [97], [93], [24],[138]
Supervised Learning [38], [15]
Semi-Supervised Learning [125]
Hybrid Learning [140]
Post-processing Statistical Evaluation [38], [125], [139], [72], [15], [140], [138]
Interpretation [31], [125], [15], [140]
Economical Evaluation [138]
Implementation [31], [80], [139], [72], [93],[140], [138]
Table 4: Processing. Papers per method cluster, based on [17].
@@
Table 5 summarises the papers according to the challenges identified in [17]. It only contains those challenges that were mentioned in the literature. The main challenge is data availability, followed by unlabeled data. Due to privacy issues, only a limited number of anonymised data is available. Additionally, since transaction data sets are often huge and resources are limited, it is hard to provide all instances with a label. This is the main obstacle for the construction of a real general benchmark. On the other hand, few papers mention feature construction, real-time execution, or verification latency as a challenge. The latter two are important when detecting transaction fraud, where decisions need to be made within seconds [118, 117]. For anti-money laundering, however, the investigation is always done after the fact, based on a longer history of transactions [70]. Hence, it is less reliant on execution time.
@@
Next to challenges in [17], we introduce four additional challenges observed in the anti-money laundering literature. We view this as an extension of the framework since these challenges can be important to fraud research in general.
@@
  • Bias in data affects the performance of the model and can be present in different ways. First, there could be bias in the missing values [79]. Second, when using synthetic data, there could be detection bias [138] present, when the synthetic data includes specific patterns that the model is trained to look for.
@@
  • Generalisation is mainly used in two ways. One way refers to the ability for methods specifically developed for anti-money laundering to be used for other fraud domains as well [140]. Another and important use of the term refers to the fact that methods need to generalise to unseen data [46], often referred to as inductive methods (compared to transductive ones).
@@
  • Anonymity of the account data gives rise to two main challenges. First, the identity of the account holder is unknown in crypto data sets, which makes it almost impossible to know if different accounts - also called wallets - belong to the same person [125]. Second, financial institution have all customer data, but due to privacy reasons, researchers cannot publish these. Hence, published articles avoid using Personally Identifiable Information all together [15].
@@
Original Automation [105], [24], [140]
Class Imbalance [125], [15]
Concept Drift [38], [15]
Data Availability [38], [23], [125], [79], [93], [138]
Feature Construction [93]
Noisy Data [105], [125], [79]
Stream Data [105], [80], [140]
Real-Time Execution [38]
Scalability [139], [93], [140]
Unlabeled Data [105], [125], [15], [140]
Verification Latency [138]
New Bias in data [79], [138]
Generalisation [140]
Robustness [72]
Anonymity [80], [125], [15]
Table 5: Challenge Papers per fraud detection challenge, based on [17].
Since this paper discusses network learning, an additional summary of the construction of the networks is given in Table 6. Most of the papers use homogeneous networks where transactions form the edges between accounts.
Paper Nodes Edges
[38] Factoring companies Economic sector geographical data transaction amount tacit link (same owners/resources)
[23] People Transactions
[105] Accounts People, businesses, accounts and location Connections
[31] People (via court register) Transactions (via bank statements)
[80] Transactions, inputs and outputs Input/output part of same transaction, inputs that belong to the same address
[125] Address Address and transactions Transactions If they are related
[79] People (from police reports) Co-appearance in the report (family, client-lawyer etc.)
[139] All items of interest (bank account, name, email, address, phone number) Link bank account to other items
[72] Bank accounts Transfers
[97] Bitcoin wallet address Transactions
[93] - -
[24] People Relations
[140] Accounts Transactions
[138] mWallets Transactions
Table 6: Network Construction. Summary of the construction of the network.
Next, we summarise the trends observed in the literature on network analytics for anti-money laundering. Here, the papers are split into three broad categories; unsupervised learning, (semi-)supervised learning and visualisation methods.
Unsupervised learning can be divided into two main streams. The first stream assigns and identifies specific roles of actors in a criminal network, based on police reports covering past criminal cases [79, 31]. The second stream introduces methods to track money flows through the network [139, 72, 97]. It is important to integrate these transaction chains, since money laundering is a process involving multiple steps and actors [70]. Next to these two main streams, the remaining paper [93] introduces a new covertness measure.
The above methods often rely on the application of centrality metrics, e.g. degree and betweenness centrality, which are combined and interpreted in different ways, leading to tailor-made solutions for specific sub-problems of anti-money laundering methods. Although detecting money laundering is in principle trying to find anomalous behaviour, very few studies deal with (unsupervised) anomaly detection [77, 99, 32].
(Semi-)Supervised learning methods deal with different aspects of money-laundering. Their strength often lies in a combination of (1) the methods used and (2) how the network is constructed, tailored to the problem at hand. The obtained network features are put into downstream machine learning tasks to classify clients/transactions. However, to the best of our knowledge. Promising methods are often based on deep learning [140], but these are mostly black boxes. To maintain interpretability, network metrics are recalculated for multiple networks, each representing a specific aspect of the data, to extract predictive features [38, 125]. However, research on the interpretability of deep network representation methods, as mentioned in [69], is still scarce.
@@ 3.4 Further Considerations
We have seen that the network analytics often relies on simple centrality measures [31, 93, 44]. The application of more complex network representation methods [117, 20] has not found mainstream adoption yet. Hence, for the construction of the experiment, we do an additional study to identify popular methods in the network analytics literature. This allows us to apply state-of-the-art network methods from other domains. On the other hand, it allows us to analyse which methods work well for money laundering and study why others might be less suited.
In this paper, the experiments are constructed for supervised learning methods. Supervised methods can easily be compared, since they can be applied to any labelled network data set. Most supervised learning methods in the literature employ standard network metrics as features. These will form the basis of our experiment, and will be supplemented with more advanced deep learning methods. The experiment will not include unsupervised learning. These are often tailor made for specific problems in money laundering, e.g., assigning roles or finding specific flow patterns. Therefore, they are not easily comparable. Further research can construct a framework to compare unsupervised learning methods on networks. However, this is outside the scope of this work.
A major challenge identified is the availability of data. Thus, a key objective of this work is to provide a comprehensive set-up incorporating multiple state-of-the-art methods. In the next section, we present our set-up and an example using an open-source data set. The accompanying code enables researchers to quickly run experiments on their own data with minimal adjustments, promoting uniform analysis across different data sets. We hope this framework will support and expand future research.
@@ 4 Experimental Set-Up
In the previous section, we reported that most papers use centrality measures in their experiments, while only few incorporate recent advances in deep learning. To the best of our knowledge, a benchmark study comparing the state-of-the-art network learning methods for anti-money laundering is still missing. This gap motivates us to carry out an extensive experimental evaluation of the most prominent methods for AML. This section describes the experimental framework that is used, including data (Section 4.2), model specification (Section 4.1), hyperparameter tuning (Section 4.3) and performance metrics (Section 4.4).

### v3

@@ 3.3 Top Cited Papers
A final step in our discussion is an in-depth analysis of the top-cited papers. These papers are well received by the scientific community and are likely of high quality. We select the $10\%$ most cited papers per year, extended with the $10\%$ most cited papers overall (see Section 2.1). This results in 20 papers as listed in Table 4. The analysis is done in two stages to gain deeper insights. First, we revisit the different categories as presented in Table 1. Second, the papers are classified according to the processing steps and challenges, as introduced by Bockel-Rickermann et al. [14].
Table 4: Summary of the top-cited papers, with indication of whether it is a review and/or crypto-related.
Reference Title
@@
Ngai et al. [100] The application of data mining techniques in financial fraud detection: A classification framework and an academic review of literature 2011 1307
@@
Fronzetti Colladon and Remondi [41] Using social network analysis to prevent money laundering 2017 200
@@
Chang et al. [22] WireVis: Visualization of Categorical, Time-Varying Data From Financial Transactions 2007 171
@@
Senator et al. [116] Financial Crimes Enforcement Network AI System (FAIS) Identifying Potential Money Laundering from Reports of Large Cash Transactions 1995 152
@@
Dreżewski et al. [33] The application of social network analysis algorithms in a system supporting money laundering detection 2015 141
McGinn et al. [89] Visualizing Dynamic Bitcoin Transaction Patterns 2016 122
@@
Gao and Ye [42] A framework for data mining-based anti-money laundering research 2007 106
@@
Wu et al. [142] Detecting Mixing Services via Mining Bitcoin Transaction Network With Hybrid Motifs 2021 94
@@
Malm and Bichler [88] Using friends for money: the positional importance of money-launderers in organized crime 2013 66
Zhou et al. [156] A Local Algorithm for Structure-Preserving Graph Cut 2017 63
Li et al. [80] FlowScope: Spotting Money Laundering Based on Graphs 2020 54
@@
Irwin and Turner [64] Illicit Bitcoin transactions: challenges in getting to the who, what, when and where 2018 47
@@
Phetsouvanh et al. [106] EGRET: Extortion Graph Exploration Techniques in the Bitcoin Network 2018 38
@@
Kute et al. [75] Deep Learning and Explainable Artificial Intelligence Techniques Applied for Detecting Money Laundering–A Critical Review 2021 38
@@
Zhdanova et al. [155] No Smurfs: Revealing Fraud Chains in Mobile Money Transfers 2014 32
@@
Bhalerao et al. [13] Mapping the underground: Supervised discovery of cybercrime supply chains 2019 24
Ovelgönne et al. [102] Covertness Centrality in Networks 2012 20
@@
Cheong and Si [23] Event-based approach to money laundering data analysis and visualization 2010 12
@@
Gerbrands et al. [47] The effect of anti-money laundering policies: an empirical network analysis 2022 11
@@
Zhou et al. [157] Behavior-Aware Account De-Anonymization on Ethereum Interaction Graph 2022 7
3.3.1 General Classification.
The results show that five out of 11 review papers are selected. This is to be expected, since review papers have a higher likelihood of being cited [32]. Plots summarising the classification according to Table 1 are given in Appendix D. Similar results are observed for the top-cited papers as for the literature as a whole; unsupervised learning is by far the most applied method, and the research relies heavily on manual and rule-based features. For the evaluation of the methods, only time and threshold-dependent metrics were used. This indicates that threshold-independent metrics are not fully mainstream when comparing AML models using networks analytics. We observe that, compared to the literature as a whole, flow/chain detection methods are better cited as well as studies using open-source data. As mentioned before, flow detection better captures the intricate patterns used by money launderers. The use of open-source data makes comparison with and replication of the studies possible, fostering adoption of the presenting research.
@@
Additionally, five papers deal with crypto-related money laundering, indicating a strong research interest. The literature mentions two causes explaining the growing interest. On the one hand, crypto currencies are widely used by criminals because of the absence of regulation [40] and the anonymity it provides [64, 139]. On the other hand, a full history of ownership is available on the blockchain for each coin, facilitating the discovery of transaction patterns in payment networks [64, 139].
Since this paper discusses network learning, an additional summary of the construction of the networks is given in Table 5. Most of the papers use homogeneous networks where transactions form the edges between accounts. Only a few use heterogeneous networks, either by having transactions as nodes as well, or having nodes that represent different characteristics or feature values.
Table 5: Summary of the construction of the network.
Paper Nodes Edges
Fronzetti Colladon and Remondi [41] Factoring companies Economic sector geographical data transaction amount tacit link (same owners/resources)
Chang et al. [22] People Transactions
Senator et al. [116] Accounts People, businesses, accounts and location Connections
Dreżewski et al. [33] People (via court register) Transactions (via bank statements)
McGinn et al. [89] Transactions, inputs and outputs Input/output part of same transaction, inputs that belong to the same address
Wu et al. [142] Address Address and transactions Transactions If they are related
Malm and Bichler [88] People (from police reports) Co-appearance in the report (family, client-lawyer etc.)
Zhou et al. [156] All items of interest (bank account, name, email, address, phone number) Link bank account to other items
Li et al. [80] Bank accounts Transfers
Phetsouvanh et al. [106] Bitcoin wallet address Transactions
Ovelgönne et al. [102] - -
Cheong and Si [23] People Relations
Zhou et al. [157] Accounts Transactions
Zhdanova et al. [155] mWallets Transactions
3.3.2 Processing steps.
@@
Next, these papers are classified according to the framework presented by Bockel-Rickermann et al. [14]. Table 6 summarises the papers over the different processing steps. The most prevalent pre-processing steps are feature engineering and exploration. The feature engineering is done to define different risk profiles or to enhance GNN models. We see in the literature that this is either based on neighbourhood and centrality metrics [41, 33, 102] or on the aggregation of local (node-specific) information [157, 155]. For exploration, some of the research is done on how to construct meaningful visualisations [22, 89], while other papers rely on clustering the nodes as additional support for their findings [88, 41].
@@
For the post-processing step, most research covers statistical evaluation, implementation and interpretation. The statistical evaluation is used to have a quantitative evaluation of the performance of the method. All research on the implementation looks at the efficiency of the methods. It deals solely with computation time and scalability. To have interpretations in the post-processing step, researchers use specific interpretable features.
Table 6: Papers per method cluster, based on Bockel-Rickermann et al. [14].
@@
Pre-processing Sampling [41].
Exploration [41],[22],[89], [88].
Missing Value Treatment [41].
@@
Categorization, standardization & segmentation [33], [13].
Feature Engineering [41], [33], [102],[157], [155].
@@
Processing Unsupervised Learning [22], [116], [33], [88], [156], [80], [106], [102], [23],[155].
Supervised Learning [41], [13].
Semi-Supervised Learning [142].
Hybrid Learning [157].
Post-processing Statistical Evaluation [41], [142], [156], [80], [13], [157], [155].
Interpretation [33], [142], [13], [157].
Economical Evaluation [155].
Implementation [33], [89], [156], [80], [102],[157], [155].
More detailed trends in the literature are discussed below according to three main categories, that mainly follow the processing steps; unsupervised learning, (semi-)supervised learning and visualisation methods.
Unsupervised learning can be divided into two main streams. The first stream assigns and identifies specific roles of actors in criminal networks, based on police reports covering past criminal cases [88, 33]. The second stream introduces methods to track money flows through the network [156, 80, 106]. It is important to integrate these transaction chains, since money laundering is a process involving multiple steps and actors [77]. Next to these two main streams, the remaining paper [102] introduces a new covertness measure.
The above-mentioned methods often rely on the application of centrality metrics, e.g. degree and betweenness centrality, which are combined and interpreted in different ways, leading to tailor-made solutions for specific sub-problems of anti-money laundering methods. Although detecting money laundering is in principle trying to find anomalous behaviour, very few studies deal with (unsupervised) anomaly detection [86, 109, 34], leaving many possibilities for future work.
(Semi-)Supervised learning methods deal with different aspects of money laundering. Their strength often lies in a combination of (1) the methods used and (2) tailoring the network to the problem at hand. The obtained network features are put into downstream classifier. To maintain interpretability, network metrics are recalculated for multiple networks, each representing a specific aspect of the data, to extract predictive features [41, 142]. Promising methods are often based on deep learning [157], but these are mostly black boxes. However, research on the interpretability of deep network representation methods, as mentioned by Kute et al. [75], is still scarce.
3.3.3 Challenges.
@@
Table 7 summarises the papers according to the challenges identified by Bockel-Rickermann et al. [14]. It only contains those challenges that were mentioned in the literature. The main challenge is data availability, followed by unlabelled data. Due to privacy issues, only a limited number of anonymised data is available. Additionally, since transaction data sets are often huge and resources are limited, it is hard to provide all instances with a label. This is the main obstacle for the construction of a general benchmark. On the other hand, few papers mention feature construction, real-time execution, or verification latency as a challenge. The latter two are important when detecting transaction fraud, where decisions need to be made within seconds [133, 131]. For anti-money laundering, however, the investigation is always done after the fact, based on a longer history of transactions [77]. Hence, it is less reliant on execution time.
@@
We introduce four additional challenges observed in the anti-money laundering literature, supplementing the work of Bockel-Rickermann et al. [14]. We view this as an extension of the framework since these challenges can be important to fraud research in general.
@@
  • Bias in data affects the performance of the model and manifests itself in different ways. First, there is bias in the missing values [88]. Second, when using synthetic data, there could be detection bias [155] present, when the synthetic data includes specific patterns that the model is trained to look for.
@@
  • Generalisation is mainly used in two ways. One way refers to the ability for methods specifically developed for anti-money laundering to be used for other fraud domains as well [157]. Another and important use of the term refers to the need for methods to generalise to unseen data [49], often referred to as inductive methods (compared to transductive ones). This is an important consideration when selecting network-based methods [55, 131, 132].
@@
  • Anonymity of the account data gives rise to two main challenges. First, the identity of the account holder is unknown in crypto data sets, which makes it almost impossible to know if different accounts - also called wallets - belong to the same person [142]. Second, financial institution possess customer data, but due to privacy reasons, researchers cannot publish these. Hence, published articles avoid using Personally Identifiable Information all together [13].
Table 7: Papers per fraud detection challenge, based on Bockel-Rickermann et al. [14].
@@
Original Automation [116], [23], [157].
Class Imbalance [142], [13].
Concept Drift [41], [13].
Data Availability [41], [22], [142], [88], [102], [155].
Feature Construction [102].
Noisy Data [116], [142], [88].
Stream Data [116], [89], [157].
Real-Time Execution [41].
Scalability [156], [102], [157].
Unlabelled Data [116], [142], [13], [157].
Verification Latency [155].
New Bias in data [88], [155].
Generalisation [157].
Robustness [80].
Anonymity [89], [142], [13].
@@ 3.4 Further Considerations
Based on the observations above, we give an intermediate conclusion and perspective on future work. The methods in the literature as a whole still rely heavily on expert-based methods, i.e., manual feature engineering and rule-based systems, while deep learning methods both for network representation as for down-stream classifiers are gaining traction. A more detailed analysis of these methods is needed to evaluate their performance properly, especially given that the current literature barely applies threshold-independent metrics meant for highly skewed data, like the AUC-PR. Moreover, our analysis showed that none of the top-cited papers apply threshold-independent metrics. This detailed analysis is what we are going to do in the second part of this paper.
Different gaps in the literature still exist which we believe gives ample opportunities for future work. A first gap was also identified by Kute et al. [75], namely the lack of interpretation tools. Future work can focus on the different interpretation methods that already exist to try and explain the output of graph neural networks.
A second gap comes from the fact that classic unsupervised methods [21, 3] do not seem to be well represented. Future work should analyse the effectiveness of unsupervised learning in networks to find anomalous behaviour. This can be beneficial since many clients and transactions are effectively unlabelled, and the modi operandi for money laundering keeps evolving. The second part of this paper will introduce a first way of applying such unsupervised learning methods, but a more extensive study is needed.
A third is that most methods are based on the analysis of static networks. Some research is being done to incorporate the time-dynamic nature of transaction networks using recurrent neural networks, but we believe that much more can still be done here.
A fourth gap is related to the inherent incompleteness of labels in AML. As criminals try to cover their tracks and financial institutions have only limited resources [133, 107], it is a given that there are still many money laundering cases that stay unnoticed. These transactions are seen as having label 0 when training the models, resembling more a positive and unlabelled (PU) problem. Specific PU-learning methods for networks should be applied when training a model.
A fifth gap is related to the limited resources as well. Since not all clients and transactions can be investigated, the predictions should take this into account. One way of dealing with this is learning-to-rank [28, 135], which has already been shown to be effective in fraud detection. The main benefit is that we focus more on correctly ranking the top most suspicious clients, while less importance is given to have a correct ranking for less suspicious clients. Although this is highly relevant, it has not yet been applied to the AML literature using network analytics.
@@ 4 Experimental Set-Up
Most papers discussed above use centrality measures, while only few incorporate deep learning. To the best of our knowledge, a benchmark study comparing the state-of-the-art network learning methods for anti-money laundering is still missing. This section describes the experimental framework to fill this gap, including model specification (Section 4.1), data (Section 4.2), hyperparameter tuning (Section 4.3) and performance metrics (Section 4.4). The focus is on supervised learning, given that the data sets used contain ground-truth labels. While some unsupervised methods are explored, a comprehensive experiment is left for future work. Visualization methods are excluded as they serve investigative support rather than objective classification. Given challenges in open-source data availability, we present results on two datasets and provide code to facilitate reproducible experiments. We hope this framework will support and expand future research, and allow uniform publishing and comparing of results on proprietary data without the need to disclose sensitive information.